9.The electric motor of a model train accelerates the train from rest to 0.620m/s in 21s. The 

total mass of the train is 875g. Find the average power delivered to the train during the 

acceleration? 


1
Expert's answer
2021-11-29T19:08:32-0500

The average power is given as follows (see https://www.khanacademy.org/science/physics/work-and-energy/work-and-energy-tutorial/a/what-is-power):


"P = \dfrac{mv_{final}^2}{2t}"

where "m = 875g = 0.875kg, v_{final} = 0.620m\/s,t = 21s". Thus, obtain:


"P = \dfrac{0.875\cdot 0.62^2}{2\cdot 21}\approx 0.008W"

Answer. 0.008W.